OpenStudy (anonymous):

Solve 4^(2x) = 7^(x−1) how do you start this problem?

OpenStudy (photon336):

\[4^{2x} = 7^{x-1}\] right?

OpenStudy (anonymous):

yes thats it

OpenStudy (photon336):

You can use logs to solve this

OpenStudy (anonymous):

oh okay, ive missed a few days of school though aand im behind how would i enter in into my calculator?

OpenStudy (anonymous):

a.2.35389 b.−2.35389 c.−1 d.1 these are the final answer options i have

OpenStudy (photon336):

okay so see \[4^{2x} = 2x*\log(4)\]

OpenStudy (anonymous):

okay so would that me the other side would equal \[7^{x-1}=x-1*\log_{7} \]

OpenStudy (anonymous):

?

OpenStudy (photon336):

so on the other side it becomes Yep :) \[2xlog(4) = \log(7)(x-1)\]

OpenStudy (photon336):

\[2x*\log(4) = x*\log(7) - \log(7)\]

OpenStudy (anonymous):

then you enter the whole equation into the calculator?

OpenStudy (photon336):

well before that we've got to get x by itself

OpenStudy (anonymous):

oh okay, so to do that we take away one x from the right side?

OpenStudy (photon336):

\[2x*\log(4)-x \log(7) = -\log(7)\] factor out a x \[x(2\log(4)-\log(7) = -\log(7)\] then divide that whole thing in the ( ) \[\frac{ -\log(7) }{ (2\log(4)-\log(7)) } = x \]

OpenStudy (anonymous):

oh okay so when you take the x away from the right side you have to take the log and everything with it then simplify it

OpenStudy (anonymous):

so then that would be what we enter into the calculator because x is by itself correct?

OpenStudy (photon336):

yeah that's the simplified form you can enter that into a calculator and get the exact value of x

OpenStudy (anonymous):

okay so i got -2.35388 which issss

OpenStudy (anonymous):

b!? right?

OpenStudy (photon336):

we can easily check this answer by plugging it into the original equation

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

that worked for me :) ty so much!

OpenStudy (photon336):

np anytime :)